# Temperature profiles of accretion disks in luminous active galactic nuclei derived from ultraviolet spectroscopic variability

### v1

@@
The characteristic timescale ($\tau$) of continuum variability of the accretion disk in active galactic nuclei (AGNs) is known to be related to the thermal timescale, which is predicted to scale with AGN luminosity ($L$) and restframe wavelength ($\lambda_{\rm RF}$) as $t_{\rm th}\propto L^{0.5}\lambda_{\rm RF}^{2}$ in the standard disk model. Using multi-epoch spectroscopic data from the Sloan Digital Sky Survey Reverberation Mapping project, we construct ultraviolet ensemble structure functions of luminous AGNs as a function of their luminosity and wavelength. Assuming that AGNs exhibit a single universal structure function when $\Delta t$ is normalized by $\tau$, wherein $\tau\propto L^{a}\lambda_{\rm RF}^{b}$, we find $a=0.50\pm 0.03$ and $b=1.42\pm 0.09$. While the value of $a$ aligns with the prediction from the standard disk model, $b$ is significantly smaller than expected, suggesting that the radial temperature (color) profile of the accretion disk is significantly steeper (shallower) than the standard disk model. Notably, this discrepancy with theory has been observed in previous studies based on spectroscopic reverberation mapping and gravitational microlensing. Although no current model of accretion disks fully matches our results, our findings provide valuable constraints for testing future physical models.
@@ 1 Introduction
While supermassive black holes (BHs) are ubiquitous in massive galaxies, only a minority of them are accreting sufficiently to shine as active galactic nuclei (AGNs). AGNs emit strong radiation over a wide range of wavelengths from the radio to the X-rays. In general, the emission from AGNs exhibits substantial variability with a timescale of days to years across all energies. In addition, variability originating from different substructures is found to be correlated (Matthews & Sandage 1963; Wagner & Witzel 1995; Ulrich et al. 1997). Under the assumption that the intrinsic variation in the accretion disk is echoed by extended structures such as the broad-line region (BLR) and the dusty torus with a time delay corresponding to the light crossing time, the radial size of such structures can be estimated by measuring time lags. Reverberation mapping (RM; Bahcall et al. 1972; Blandford & McKee 1982), first applied to probe the size of the BLR using the time delay between the variation in the accretion disk continuum and the broad emission lines (e.g., Peterson et al. 1994; O’Brien et al. 1998), subsequently led to the development of methods for estimating the masses of BHs in AGNs (e.g., Gebhardt et al. 2000; Peterson et al. 2004; Greene & Ho 2005; Vestergaard & Peterson 2006; Ho & Kim 2015; Wang & Woo 2024).
@@
The RM method also provides a powerful probe of the size of the accretion disk (Krolik et al. 1991; Wanders et al. 1997). Simplifying the structure of an accretion disk into a series of annuli with a negative temperature gradient (i.e., lower temperature at larger radii), the time lag between the accretion disk continuum at different wavelengths, originating from different temperatures, can serve as a proxy for the disk size. Intriguingly, disk sizes estimated from continuum RM are systematically larger (e.g., Fausnaugh et al. 2016; but see Homayouni et al. 2019) than predicted from the standard disk model (Shakura & Sunyaev 1973; Frank et al. 2002). A similar trend has also been identified through gravitational microlensing studies (e.g., Morgan et al. 2010). Various explanations have surfaced to explain this “disk size problem,” most based on a modification of the geometry or emitting source: accretion disk blown up by a disk wind (Sun et al. 2019), rimmed or rippled accretion disk (Starkey et al. 2023), X-ray source with enhanced height (Kammoun et al. 2019), disk reprocessing in the X-rays or far-ultraviolet (Gardner & Done 2017), or a significant contribution from the recombination continuum of the BLR (e.g., McHardy et al. 2018).
@@
The radial profile of the effective temperature, defined as $T\propto R^{-1/p}$, can be used to test the validity of accretion disk models, as different models predict different temperature profiles. From an observational point of view, under the assumption that the radiation follows a blackbody spectrum ($\lambda_{\rm max}\propto T^{-1}$), it is more common to measure the more observationally tractable radial color profile, $R\propto\lambda_{\rm RF}^{p}$, where $\lambda_{\rm RF}$ is the restframe wavelength. For example, while $p=4/3$ is expected from the standard disk model, a larger value of $p$ is anticipated from the disk wind model (Li et al. 2019; Sun et al. 2019), while the opposite holds for the magneto-rotational instability model (Agol & Krolik 2000). Continuum RM studies find that the observed time lags for various wavelengths are well fit with $p\approx 4/3$ as expected from the standard disk model (McHardy et al. 2014; Homayouni et al. 2019). However, both steeper ($p<4/3$; Fausnaugh et al. 2016; Starkey et al. 2017) and shallower ($p>4/3$; Jiang et al. 2017) profiles have been detected. The results from microlensing studies are more controversial, reporting a wide range of values for $p$. While many claim that $p$ is consistent with $4/3$ (e.g., Eigenbrod et al. 2008; Poindexter et al. 2008; Rivera et al. 2023), values of $p$ significantly less than $4/3$ have been derived by Floyd et al. (2009), Blackburne et al. (2011), Jiménez-Vicente et al. (2014), and Muñoz et al. (2016). By contrast, Bate et al. (2018) and Cornachione & Morgan (2020) note an opposite trend. The origin of these discrepant results is unclear. Certainly, small-number statistics come into play, and it is difficult to judge the impact of sample selection and experimental methodology. It is vital to investigate the radial color profile for a large sample of AGNs using an independent method.
@@
As a complementary approach, color profiles can be quantified using the characteristics of ultraviolet-optical variability arising from the accretion disk. This intrinsic variability can be characterized using the power spectral density, defined as the variability power as a function of frequency $\nu$. One can also compute the structure function (SF), defined as the variability amplitude as a function of time lag $\Delta t$. The power spectral density and SF of the accretion disk continuum are well fit with a broken power-law in the sense that the variability amplitude increases with increasing timescale and flattens above a certain characteristic timescale ($\tau$; Kelly et al. 2009; MacLeod et al. 2010; Guo et al. 2017). Previous studies showed that $\tau$ corresponds to the thermal timescale [^1] The thermal timescale is the approximate time it takes to establish thermal equilibrium. $t_{\rm th}$ (Kelly et al. 2009; Sun et al. 2015; Burke et al. 2021; Arévalo et al. 2024; but see MacLeod et al. 2010). Although most studies reached a similar conclusion, their relations between $\tau$ and AGN properties vary, possibly due to systematic uncertainties in the measurements of $\tau$ caused by imperfect light curves with a sparse cadence and a relatively short baseline (Emmanoulopoulos et al. 2010; Kozlowski 2017).
@@
Alternatively, the dependence of $\tau$ on the AGN luminosity and wavelength can be investigated by assuming that all the SFs or power spectral densities follow a single universal function if $\Delta t$ or $\nu$ is converted to the units of the characteristic timescale. Applying this method to the broadband light curves of thousands of AGNs, Tang et al. (2023) demonstrated that $\tau\propto L^{0.539\pm 0.004}\lambda_{\rm RF,broad}^{2.418\pm 0.023}$, where $\lambda_{\rm RF}$ is the effective wavelength of the broadband filter in the restframe. Although the slopes of the luminosity and wavelength factor are marginally steeper than expected from the standard disk model, this finding is broadly consistent with the predictions of the standard disk model on the thermal timescale (i.e., $\tau\propto L^{0.5}\lambda_{\rm RF}^{2}$; but see Arévalo et al. 2024; Petrecca et al. 2024). Note that, in the standard disk model, the thermal timescale is expected to correlate with AGN luminosity and wavelength as $\tau\propto L^{0.5}\lambda_{\rm RF}^{2}$ (Shakura & Sunyaev 1973; Frank et al. 2002). This discrepancy is possibly due to several shortcomings of broadband photometry, including (1) contamination from broad emission lines and iron multiples (e.g., Boroson & Green 1992), (2) the broad range of the effective wavelength in the restframe depending on the redshift of the targets (e.g., MacLeod et al. 2010), and (3) time variation of the shape of the continuum (e.g., Guo & Gu 2016), which can introduce possible bias in converting broadband photometry to continuum flux. Despite these limitations, broadband monitoring data has been widely used, for they can be obtained efficiently for large samples compared to more resource-intensive spectroscopic monitoring, which is available only for more limited samples.
@@
To properly investigate the dependence of $\tau$ on the restframe wavelength, with the aim of constraining the temperature profile of the accretion disk, it is preferable to measure the continuum variability using spectroscopic data, which can mitigate more effectively contaminants such as strong emission lines. The Sloan Digital Sky Survey Reverberation Mapping (SDSS-RM) project (Shen et al. 2015a, 2024), which has conducted intensive spectroscopic monitoring of a relatively large sample of quasars, provides an opportunity to carry out this experiment. With this dataset, we attempt to examine the dependence of the characteristic timescale of continuum variability on AGN luminosity and wavelength using spectroscopically determined continuum light curves.
@@
Figure 1: Systematic uncertainty in magnitude due to uncertainty in flux calibration ($\sigma_{\rm sys}$) as a function of the observed wavelength and flux density. The systematic uncertainties increase dramatically at the blue and red ends of the spectra ($\lambda_{\rm obs}<5000\,{\rm\AA}$ and $\lambda_{\rm obs}>9000\,{\rm\AA}$)
@@ 2 Sample and data
The continuum variability of AGNs typically is studied using broadband photometry (e.g., Vanden Berk et al. 2004; MacLeod et al. 2010; Caplar et al. 2017; Tang et al. 2023; Arévalo et al. 2024). A shortcoming of this conventional method is that it can be contaminated easily by strong emissions from the BLR, including broad emission lines, the myriad blends from iron multiplets, and the Balmer continuum. By comparison, spectroscopic data offers the advantage that specific wavelength windows can be chosen to reject strong contaminants. This is the strategy we adopt in this study. We utilize the spectroscopic monitoring data from the SDSS-RM project [^2] https://ariel.astro.illinois.edu/sdssrm/final_result/ (Shen et al. 2015a, 2024) to construct the ultraviolet continuum SF. The SDSS-RM project monitored a flux-limited ($i_{\rm PSF}<21.7$ mag) sample of 849 quasars at $4.50.1<z<$ 4.5 using a 2.5 m SDSS telescope with the BOSS spectrograph, covering the wavelength range $3650-10400\,\rm{\AA}$ at a spectral resolution of $\sim 2000$. Light curves from this project span $\sim 7$ years with an average cadence of $\sim 4$, 12, and 30 days for the 2014, 2015--2017, and 2018--2020 observing seasons, respectively. Because accurate measurement of the continuum fluxes and their uncertainties is crucial to estimating the SF robustly, we compute those values directly from the spectra instead of extracting them from the best-fit continuum, which may lead to systematic bias.
@@ 2.1 Measurements of flux and its uncertainty
To measure the continuum variability using spectroscopic data, we define continuum windows at various wavelengths and estimate the flux densities from the single-epoch spectra within these regions. The uncertainties of the flux density are given in each spectral pixel for the single-epoch spectra. However, additional uncertainties stemming from imperfect flux calibration (e.g., Margala et al. 2016; Milaković et al. 2021) should be taken into account. We estimate the final uncertainty of the continuum flux ($\sigma_{\rm final}$) as the square root of the quadratic sum of these two sources of uncertainty.
@@
The uncertainty measured from the continuum window with a width of 40 ${\rm\AA}$ is calculated as $\sigma_{\rm flux}/\sqrt{N}$, where $\sigma_{\rm flux}$ is half the 16th to 84th percentile range of the flux density, and $N$ is a number of data points within the continuum window. We use a median of the flux density within the continuum window as the flux at the continuum. For the systematic uncertainty due to the unstable flux calibration, we statistically estimated it using the multi-epoch spectra of our sample. If we assume that our sample does not vary within 2 days (i.e., ${\rm SF_{2\,days}\approx 0}$), the observed flux variation within this time interval is caused solely by the systematics in the flux calibration. This systematic uncertainty, in units of magnitude, can be expressed as $\sigma_{\rm sys}=\sqrt{(\sigma_{\rm\Delta mag}/\sqrt{2})^{2}-\sigma_{\rm mag}^{2}}$, where $\sigma_{\rm\Delta mag}$ is half the 16th to 84th percentile range of the differences between magnitude pairs observed within 2 days [^3] Note that the pair within 2 days is selected in the observed frame. Therefore, in the restframe, $\Delta t$ between the pair is on average $\sim 0.76\pm 0.21$ days., and $\sigma_{\rm mag}$ is the uncertainty of the continuum flux in the single-epoch spectrum in units of magnitude (Figure 1). This value cannot be estimated for an individual object because the spectroscopic pairs observed within $2$ days are typically fewer than seven for a single target, which is not sufficient to estimate the uncertainty. Instead, we derive the systematic uncertainties using the entire sample as a function of the observed wavelength ($\lambda_{\rm obs}$) and flux density ($F_{\lambda}$). We find that it is sensitive to the observed wavelength and the brightness of the target (Figure 1). The systematic uncertainty of the single spectra at a specific wavelength is determined by interpolation.
@@
Intriguingly, $\sigma_{\rm sys}$ is substantial at the spectral edges (e.g., $\lambda_{\rm obs}<5000\,{\rm\AA}$ and $\lambda_{\rm obs}>9000{\rm\AA}$), indicating that the relative flux calibration error and imperfect sky subtraction may be severe in the original data. Motivated by this result, we opt not to use the spectral data at $\lambda_{\rm obs}<5000\,{\rm\AA}$ or $\lambda_{\rm obs}>9000{\rm\AA}$ in the observed frame. Additionally, we do not use objects with $\sigma_{\rm final}\geq 0.11$ mag, for which the noise dominates the variability because we cannot measure the error exactly for each object, but only statistically. Finally, to investigate the variability solely from the accretion disk, we attempt to avoid the spectral regions that can be highly contaminated by broad emission lines, the pseudo-continuum originating from Fe II multiplets (Shen et al. 2011), and the stellar continuum of the host galaxy. Consequently, the continuum fluxes are measured as median values within 40 ${\rm\AA}$-wide continuum windows centered at 1367, 1479, 1746, 2230, 3050, 3500, and 3800 ${\rm\AA}$ in the rest frame. Note that the continuum fluxes at 3050 and 3500 Å may be influenced by the Balmer continuum (e.g., Cackett et al. 2018; Korista & Goad 2019). This phenomenon is further explored in Section 5.1. The width of the continuum window is set empirically to be wide enough to secure high signal-to-noise ratio fluxes (S/N $\geq 3$) [^4] More than 99% of the continuum measurements for the variable targets have an S/N of 3 or higher. while at the same time minimizing the contribution from contaminants.
@@
Figure 2: Distribution of BH mass and bolometric luminosity of the final sample with available BH mass measurements. The points are color-coded by redshift. The dotted lines represent Eddington ratios of $\lambda_{\rm Edd}=0.01$, $0.1$, and $1$.

### v2

@@
The characteristic timescale ($\tau$) of the continuum variability of the accretion disk in active galactic nuclei (AGNs) is known to be related to the thermal timescale, which is predicted to scale with AGN luminosity ($L$) and the rest-frame wavelength ($\lambda_{\rm RF}$) as $t_{\rm th}\propto L^{0.5}\lambda_{\rm RF}^{2}$ in the standard disk model. Using multi-epoch spectroscopic data from the Sloan Digital Sky Survey Reverberation Mapping project, we constructed ultraviolet ensemble structure functions of luminous AGNs as a function of their luminosity and wavelength. Assuming that AGNs exhibit a single universal structure function when $\Delta t$ is normalized by $\tau$, wherein $\tau\propto L^{a}\lambda_{\rm RF}^{b}$, we find $a=0.50\pm 0.03$ and $b=1.42\pm 0.09$. While the value of $a$ aligns with the prediction from the standard disk model, $b$ is significantly smaller than expected, suggesting that the radial temperature (color) profile of the accretion disk is significantly steeper (shallower) than the standard disk model. Notably, this discrepancy with theory has been observed in previous studies based on spectroscopic reverberation mapping and gravitational microlensing. Although no current model of accretion disks fully matches our results, our findings provide valuable constraints for testing future physical models.
@@ 1 Introduction
While supermassive black holes (BHs) are ubiquitous in massive galaxies, only a minority of them accrete sufficiently to shine as active galactic nuclei (AGNs). AGNs emit strong radiation over a wide range of wavelengths, from the radio to X-rays. In general, the emission from AGNs exhibits substantial variability with a timescale of days to years across all energies. In addition, variability originating from different substructures is found to be correlated (Matthews & Sandage 1963; Wagner & Witzel 1995; Ulrich et al. 1997). Under the assumption that the intrinsic variation in the accretion disk is echoed by extended structures such as the broad-line region (BLR) and the dusty torus with a time delay corresponding to the light crossing time, the radial size of such structures can be estimated by measuring time lags. Reverberation mapping (RM; Bahcall et al. 1972; Blandford & McKee 1982), first applied to probe the size of the BLR using the time delay between the variation in the accretion disk continuum and the broad emission lines (e.g., Peterson et al. 1994; O’Brien et al. 1998), led to the development of methods for estimating the masses of BHs in AGNs (e.g., Gebhardt et al. 2000; Peterson et al. 2004; Greene & Ho 2005; Vestergaard & Peterson 2006; Ho & Kim 2015; Wang & Woo 2024).
@@
The RM method also provides a powerful probe of the size of the accretion disk (Krolik et al. 1991; Wanders et al. 1997). Simplifying the structure of an accretion disk into a series of annuli with a negative temperature gradient (i.e., lower temperatures at larger radii), the time lag between the accretion disk continuum at different wavelengths, originating from different temperatures, can serve as a proxy for the disk size. Intriguingly, disk sizes estimated from continuum RM are systematically larger (e.g., Fausnaugh et al. 2016; but see Homayouni et al. 2019) than predicted from the standard disk model (Shakura & Sunyaev 1973; Frank et al. 2002). A similar trend has also been identified through gravitational microlensing studies (e.g., Morgan et al. 2010). Various explanations have surfaced to explain this “disk size problem,” most of which are based on a modification of the geometry or emitting source: an accretion disk with a disk wind (Sun et al. 2019), a rimmed or rippled accretion disk (Starkey et al. 2023), an X-ray source with an increased height (Kammoun et al. 2019), a disk reprocessing in the X-rays or far-ultraviolet (Gardner & Done 2017), or a significant contribution from the recombination continuum of the BLR (e.g., McHardy et al. 2018).
@@
The radial profile of the effective temperature, defined as $T\propto R^{-1/p}$, can be used to test the validity of accretion disk models, as different models predict different temperature profiles. From an observational point of view, under the assumption that the radiation follows a blackbody spectrum ($\lambda_{\rm max}\propto T^{-1}$), it is more common to measure the more observationally tractable radial color profile, $R\propto\lambda_{\rm RF}^{p}$, where $\lambda_{\rm RF}$ is the rest-frame wavelength. For example, while $p=4/3$ is expected from the standard disk model, a larger value of $p$ is anticipated from the disk wind model (Li et al. 2019; Sun et al. 2019), and the opposite holds for the magnetorotational instability model (Agol & Krolik 2000). Continuum RM studies find that the observed time lags for various wavelengths are well fit with $p\approx 4/3$, as expected from the standard disk model (McHardy et al. 2014; Homayouni et al. 2019). However, both steeper ($p<4/3$; Fausnaugh et al. 2016; Starkey et al. 2017) and shallower ($p>4/3$; Jiang et al. 2017) temperature profiles have been detected. The results from microlensing studies, which report a wide range of values for $p$, are more debated. While many claim that $p$ is consistent with $4/3$ (e.g., Eigenbrod et al. 2008; Poindexter et al. 2008; Rivera et al. 2023), values of $p$ lower than $4/3$ have been derived by Floyd et al. (2009), Blackburne et al. (2011), Jiménez-Vicente et al. (2014), and Muñoz et al. (2016), and values of $p$ higher than $4/3$ have been derived by Bate et al. (2018) and Cornachione & Morgan (2020). The origin of these discrepant results is unclear. Certainly, small-number statistics come into play, and it is difficult to judge the impact of sample selection and experimental methodology. It is vital to investigate the radial color profile for a large sample of AGNs using an independent method.
@@
As a complementary approach, color profiles can be quantified using the characteristics of ultraviolet-optical variability arising from the accretion disk. This intrinsic variability can be characterized using the power spectral density, defined as the variability power as a function of the frequency ($\nu$). One can also compute the structure function (SF), defined as the variability amplitude as a function of the time lag ($\Delta t$). The power spectral density and SF of the accretion disk continuum are well fit with a broken power law in the sense that the variability amplitude increases with increasing timescale and flattens above a certain characteristic timescale ($\tau$; Kelly et al. 2009; MacLeod et al. 2010; Guo et al. 2017). Studies have shown that $\tau$ corresponds to the thermal timescale [^1] The thermal timescale is the approximate time it takes to establish thermal equilibrium., $t_{\rm th}$ (Kelly et al. 2009; Sun et al. 2015; Burke et al. 2021; Arévalo et al. 2024; but see MacLeod et al. 2010). Although most studies reached a similar conclusion, the relations they found between $\tau$ and AGN properties vary, possibly due to systematic uncertainties in the measurements of $\tau$ caused by imperfect light curves with a sparse cadence and a relatively short baseline (Emmanoulopoulos et al. 2010; Kozlowski 2017).
@@
Alternatively, the dependence of $\tau$ on the AGN luminosity and wavelength can be investigated by assuming that all the SFs or power spectral densities follow a single universal function if $\Delta t$ or $\nu$ is converted to the units of the characteristic timescale. Applying this method to the broadband light curves of thousands of AGNs, Tang et al. (2023) demonstrated that $\tau\propto L^{0.539\pm 0.004}\lambda_{\rm RF,broad}^{2.418\pm 0.023}$, where $\lambda_{\rm RF}$ is the effective wavelength of the broadband filter in the rest-frame. Although the slopes of the luminosity and wavelength factor are marginally steeper than expected from the standard disk model, this finding is broadly consistent with the predictions of the standard disk model on the thermal timescale (i.e., $\tau\propto L^{0.5}\lambda_{\rm RF}^{2}$; but see Arévalo et al. 2024; Petrecca et al. 2024). It should be noted that, in the standard disk model, the thermal timescale is expected to correlate with the AGN luminosity and wavelength as $\tau\propto L^{0.5}\lambda_{\rm RF}^{2}$ (Shakura & Sunyaev 1973; Frank et al. 2002). This discrepancy is possibly due to several shortcomings of broadband photometry, including (1) contamination from broad emission lines and iron multiples (e.g., Boroson & Green 1992), (2) the broad range of the effective wavelength in the rest-frame, which depends on the redshift of the targets (e.g., MacLeod et al. 2010), and (3) time variation of the shape of the continuum (e.g., Guo & Gu 2016), which can introduce bias when converting broadband photometry to continuum flux. Despite these limitations, broadband monitoring data have been widely used, for they can be obtained for large samples more efficiently than the more resource-intensive spectroscopic monitoring, which is available only for more limited samples.
@@
To properly investigate the dependence of $\tau$ on the rest-frame wavelength, with the aim of constraining the temperature profile of the accretion disk, it is preferable to measure the continuum variability using spectroscopic data, which can more effectively mitigate contaminants such as strong emission lines. The Sloan Digital Sky Survey Reverberation Mapping (SDSS-RM) project (Shen et al. 2015a, 2024), which has conducted intensive spectroscopic monitoring of a relatively large sample of quasars, provides an opportunity to carry out this experiment. With this dataset, we examined the dependence of the characteristic timescale of continuum variability on AGN luminosity and wavelength using spectroscopically determined continuum light curves.
@@
Figure 1: Systematic uncertainty in magnitude due to the uncertainty in flux calibration ($\sigma_{\rm sys}$) as a function of the observed wavelength and flux density. The systematic uncertainties increase dramatically at the blue and red ends of the spectra ($\lambda_{\rm obs}<5000\,{\rm\AA}$ and $\lambda_{\rm obs}>9000\,{\rm\AA}$).
@@ 2 Sample and data
The continuum variability of AGNs typically is studied using broadband photometry (e.g., Vanden Berk et al. 2004; MacLeod et al. 2010; Caplar et al. 2017; Tang et al. 2023; Arévalo et al. 2024). A shortcoming of this conventional method is that broadband photometry can easily be contaminated by strong emissions from the BLR, including broad emission lines, the myriad blends from iron multiplets, and the Balmer continuum. By comparison, spectroscopic data offer the advantage that specific wavelength windows can be chosen to reject strong contaminants. This is the strategy we adopted in this study. We utilized the spectroscopic monitoring data from the SDSS-RM project [^2] https://ariel.astro.illinois.edu/sdssrm/final_result/ (Shen et al. 2015a, 2024) to construct the ultraviolet continuum SF. The SDSS-RM project monitored a flux-limited sample ($i_{\rm PSF}<21.7$ mag) of 849 quasars at $4.50.1<z<$ 4.5 using a 2.5 m SDSS telescope with the Baryon Oscillation Spectroscopic Survey (BOSS) spectrograph, covering the wavelength range $3650-10400\,\rm{\AA}$ at a spectral resolution of $\sim 2000$. Light curves from this project span $\sim 7$ years with an average cadence of $\sim 4$, 12, and 30 days for the 2014, 2015--2017, and 2018--2020 observing seasons, respectively. Because accurate measurement of the continuum fluxes and their uncertainties is crucial to estimating the SF robustly, we computed those values directly from the spectra instead of extracting them from the best-fit continuum, which can lead to systematic bias.
@@ 2.1 Measurements of flux and its uncertainty
To measure the continuum variability using spectroscopic data, we defined continuum windows at various wavelengths and estimated the flux densities from the single-epoch spectra within these regions. The uncertainties of the flux density are given in each spectral pixel for the single-epoch spectra. However, additional uncertainties stemming from imperfect flux calibration (e.g., Margala et al. 2016; Milaković et al. 2021) should be taken into account. We estimated the final uncertainty of the continuum flux ($\sigma_{\rm final}$) as the square root of the quadratic sum of these two sources of uncertainty.
@@
The uncertainty measured from the continuum window with a width of 40 ${\rm\AA}$ was calculated as $\sigma_{\rm flux}/\sqrt{N}$, where $\sigma_{\rm flux}$ is half the 16th to 84th percentile range of the flux density, and $N$ is the number of data points within the continuum window. We used the median of the flux density within the continuum window as the flux at the continuum. We statistically estimated the systematic uncertainty due to the unstable flux calibration using the multi-epoch spectra of our sample. If we assume that our sample does not vary over a period of 2 days (i.e., ${\rm SF_{2\,days}\approx 0}$), the observed flux variation within this time interval is caused solely by the systematics in the flux calibration. This systematic uncertainty, in units of magnitude, can be expressed as $\sigma_{\rm sys}=\sqrt{(\sigma_{\rm\Delta mag}/\sqrt{2})^{2}-\sigma_{\rm mag}^{2}}$, where $\sigma_{\rm\Delta mag}$ is half the 16th to 84th percentile range of the differences between magnitude pairs observed within 2 days of one another [^3] Note that the pair is selected in the observed frame. Therefore, in the rest-frame, the $\Delta t$ between the pair is on average $\sim 0.76\pm 0.21$ days., and $\sigma_{\rm mag}$ is the uncertainty of the continuum flux in the single-epoch spectrum in units of magnitude (Fig. 1). This value cannot be estimated for an individual object because there are typically fewer than seven spectroscopic pairs observed within $2$ days for a single target, which is not sufficient to estimate the uncertainty. Instead, we derived the systematic uncertainties using the entire sample as a function of the observed wavelength ($\lambda_{\rm obs}$) and flux density ($F_{\lambda}$). We find that it is sensitive to the observed wavelength and the brightness of the target (Fig. 1). The systematic uncertainty of the single spectra at a specific wavelength was determined via interpolation.
@@
Intriguingly, $\sigma_{\rm sys}$ is substantial at the spectral edges (e.g., $\lambda_{\rm obs}<5000\,{\rm\AA}$ and $\lambda_{\rm obs}>9000{\rm\AA}$), indicating that the relative flux calibration error and imperfect sky subtraction may be severe in the original data. Motivated by this result, we opted not to use the spectral data at $\lambda_{\rm obs}<5000\,{\rm\AA}$ or $\lambda_{\rm obs}>9000{\rm\AA}$ in the observed frame. Additionally, we did not use objects with $\sigma_{\rm final}\geq 0.11$ mag, for which the noise dominates the variability, because we cannot measure the error exactly for each object, we can only measure it statistically. Finally, to investigate the variability solely from the accretion disk, we attempted to avoid the spectral regions that can be highly contaminated by broad emission lines, the pseudo-continuum originating from Fe II multiplets (Shen et al. 2011), and the stellar continuum of the host galaxy. Consequently, the continuum fluxes were measured as median values within ${\rm\AA}$-wide continuum windows centered at 1367, 1479, 1746, 2230, 3050, 3500, and 3800 ${\rm\AA}$ in the rest-frame. We note that the continuum fluxes at 3050 and 3500 Å may be influenced by the Balmer continuum (e.g., Cackett et al. 2018; Korista & Goad 2019). This phenomenon is further explored in Sect. 5.1. The width of the continuum window was set empirically to be wide enough to secure high signal-to-noise ratio fluxes (S/N $\geq 3$) [^4] More than 99% of the continuum measurements for the variable targets have a S/N of 3 or higher. while at the same time minimizing the contribution from contaminants.
@@
Figure 2: Distribution of the BH mass and bolometric luminosity of the final sample with available BH mass measurements. The points are color-coded by redshift. The dotted lines mark Eddington ratios of $\lambda_{\rm Edd}=0.01$, $0.1$, and $1$.